# SAGE: Strategy-Adaptive Generation Engine for Query Rewriting

### v1

@@ Abstract
Query rewriting is pivotal for enhancing dense retrieval, yet current methods demand large-scale supervised data or suffer from inefficient reinforcement learning (RL) exploration. In this work, we first establish that guiding Large Language Models (LLMs) with a concise set of expert-crafted strategies, such as semantic expansion and entity disambiguation, substantially improves retrieval effectiveness on challenging benchmarks, including HotpotQA, FEVER, NFCorpus, and SciFact. Building on this insight, we introduce the Strategy-Adaptive Generation Engine (SAGE), which operationalizes these strategies in an RL framework. SAGE introduces two novel reward shaping mechanisms- Strategic Credit Shaping (SCS) and Contrastive Reward Shaping (CRS)-to deliver more informative learning signals. This strategy-guided approach not only achieves new state-of-the-art NDCG@10 results, but also uncovers a compelling emergent behavior: the agent learns to select optimal strategies, reduces unnecessary exploration, and generates concise rewrites, lowering inference cost without sacrificing performance. Our findings demonstrate that strategy-guided RL, enhanced with nuanced reward shaping, offers a scalable, efficient, and more interpretable paradigm for developing the next generation of robust information retrieval systems.
@@ 1 Introduction
Effective information retrieval (IR) is increasingly reliant on dense retrieval systems, which map queries and documents into a shared semantic space. The performance of these systems, however, is fundamentally bound by the quality of the input query. To bridge the significant gap between a user’s initial intent and a query optimized for machine comprehension, query rewriting has emerged as a critical component. While LLMs have shown significant promise for general data generation, comprehension, and reasoning Achiam et al. (2023); Wang et al. (2025a); Yang et al. (2025a, b); Wang et al. (2025b), current methodologies face two primary obstacles: traditional supervised fine-tuning demands large-scale, costly manual annotations, whereas modern RL approaches, such as PPO Schulman et al. (2017) and GRPO Shao et al. (2024), often struggle with inefficient exploration. This inefficiency not only hampers the discovery of optimal rewriting strategies but can also result in unstable training dynamics and even catastrophic failures, where models produce incoherent or irrelevant outputs.
@@
Figure 1: An overview of the Strategy-Adaptive Generation Engine (SAGE) framework. SAGE operationalizes expert-crafted strategies within a reinforcement learning loop. The policy model generates a strategy-guided rewrite, which is evaluated against the environment to produce an initial reward (NDCG@10). This reward is then transformed by our novel shaping modules-Strategic Credit Shaping or Contrastive Reward Shaping-to create a potent learning signal for the agent.
@@
The successful application of on-policy reinforcement learning to fine-tune LLMs is pioneered by algorithms such as PPO Schulman et al. (2017), which in turn inspired a family of variants including GRPO Shao et al. (2024), VAPO Yue et al. (2025), and DAPO Yu et al. (2025). While these powerful algorithms have enabled new capabilities in query rewriting for LLMs, many prior approaches Jiang et al. (2025); Wang et al. (2025c) simply adopt them as generic optimization tools, without incorporating domain-specific guidance or explicit strategies tailored to the query rewriting task.
@@
To bridge this gap, we introduce the Strategy-Adaptive Generation Engine (SAGE), as demonstrated in Fig. 1, a framework that directly integrates our human-designed strategies into the GRPO algorithm, steering LLMs toward more effective and improved query-rewriting policies. To further refine the learning signal, we also propose and investigate several novel reward shaping schemes. Beyond using a direct NDCG@10 score, these include Strategic Credit Shaping (SCS), which assigns credit based on the collective performance of each strategy, and Contrastive Reward Shaping(CRS), which recasts absolute scores into a measure of relative performance.
@@
Upon reproducing prior approaches Li et al. (2024); Jiang et al. (2025), we identify a significant form of reward hacking Weng (2024); Wang et al. (2024a); Wen et al. (2025) intrinsic to the query rewriting task. When fine-tuning LLMs on datasets such as HotpotQA Yang et al. (2018), the models consistently converge to a trivial policy: leaving the input query entirely unchanged. This behavior stems from the optimization objective, as modern retrievers like BGE-en-v1.5 Xiao et al. (2023) can already achieve strong baseline performance with the original queries. Given direct access to these queries, the model exploits this “safe” strategy-copying the input-to consistently secure high rewards, without engaging in the intended semantic rewriting. As a result, the agent becomes trapped in a deceptive local optimum, severely restricting exploration and preventing the discovery of potentially superior, albeit riskier, rewriting strategies.
@@
To counteract this, we explicitly promote exploration through targeted prompt engineering and introducing a penalty for outputs identical to the original query. This straightforward yet powerful mechanism encourages the agent to venture beyond the safe default, resulting in significantly enhanced retrieval performance. We further examine the effectiveness of this exploration strategy through detailed ablation studies in Section 5.3.
@@
We evaluate SAGE on two challenging benchmarks: HotpotQA Yang et al. (2018) and NFCorpus Boteva et al. (2016). Our results demonstrate that SAGE achieves state-of-the-art retrieval effectiveness as measured by NDCG@10. Notably, we observe an emergent behavior where SAGE learns a more efficient reasoning process, substantially reducing token usage. This improved efficiency directly translates to lower inference latency and reduced computational costs.
@@
    We demonstrate that explicitly guiding LLMs with a small set of interpretable human-designed strategies, such as semantic expansion and entity disambiguation, significantly improves query rewriting quality, even without additional training. This highlights the upper bound of performance achievable purely through prompting.
@@
    We introduce SAGE (Strategy-Adaptive Generation Engine), a novel reinforcement learning framework that systematically integrates these explicit strategies into the learning loop. SAGE autonomously adapts its strategy selection process using our proposed reward-shaping mechanisms, enabling effective refinement of the training procedure.
@@
    We establish a new state-of-the-art in dense retrieval effectiveness (NDCG@10) using SAGE, while identifying a notable emergent behavior: SAGE learns a more efficient reasoning process, substantially reducing inference latency and computational costs without explicit optimization.
@@
2 Related Work
2.1 The Evolution of Query Rewriting
2.2 Reinforcement Learning for Query Rewriting
To alleviate reliance on explicit supervision, recent approaches leverage RL to guide the generative capabilities of LLMs using weaker signals Schulman et al. (2017); Shao et al. (2024); Jiang et al. (2025); Li et al. (2024); Wang et al. (2025b, c, d). However, the application of these powerful algorithms often reveals significant limitations. A common approach is to treat the LLM as a black box, applying on-policy algorithms like PPO without modification Jiang et al. (2025), which overlooks the need for specialized guidance. Other lines of work focus on prompt engineering, but strategies tailored for sparse retrieval in general web search Li et al. (2024) exhibit limited effectiveness when transferred to dense retrieval tasks. Furthermore, even approaches that integrate reward models, such as Wang et al. (2025c), can be flawed; they often rely on an arbitrary, static fusion of scores from multiple fine-tuned reward models, lacking dynamic feedback from the actual retrieval environment. This highlights a fundamental gap: existing RL-based methods frequently neglect the nuanced, domain-specific guidance required to fully exploit the potential of dense retrieval systems. In contrast, our SAGE framework addresses this gap by introducing a novel layer of explicit, human-designed strategic guidance. It achieves this through two core innovations: first, by equipping the agent with a set of fine-grained strategies for query augmentation, and second, by proposing novel reward calculation schemes like SCS and CRS to provide a more nuanced and effective learning signal.
2.3 Challenge in RL: Reward Hacking
@@ 3 Methodology
The core idea of our methodology is to replace opaque black-box optimization with explicit, human-interpretable decision-making steps in the reinforcement learning process. To achieve this, we propose the Strategy-Adaptive Generation Engine (SAGE), which reformulates query rewriting from unconstrained text generation into a structured, strategy-driven decision problem. As shown in Figure 1, our framework embeds a set of expert-crafted strategies directly within the RL training process to systematically guide both exploration and policy learning.
@@ 3.1 Problem Formulation
We frame query rewriting as a reinforcement learning task. Given an initial user query $q_{orig}$ and a document collection $\mathcal{D}$, the objective is to learn a policy $\pi$ that generates an improved query $q$. The environment evaluates the rewritten query using NDCG@10, yielding an initial reward $r_{\text{orig}}=\text{NDCG@10}(q,\mathcal{D})$, which is further refined by our reward shaping methods (see Section 3.4) to produce the final reward $r_{\text{final}}$. The agent thus aims to maximize the expected final reward: $\mathbb{E}_{q\sim\pi(\cdot|q_{orig})}[r_{\text{final}}]$. We use GRPO Shao et al. (2024) as our optimization algorithm and incorporate a set of expert-crafted rewriting strategies $\mathcal{S}=\{s_{1},s_{2},\dots,s_{5}\}$, providing explicit guidance for query reformulation.
Strategy Targeted Challenge Primary Use Case
Senmantic Expansion Vocabulary mismatch between query and documents. General purpose, especially in specialized domains (e.g., NFCorpus).
Entity Disambiguation Ambiguous entities leading to incorrect retrieval. Queries with common names or acronyms.
Sub-question Decomposition Complex, multi-step information needs. Multi-hop QA (e.g., HotpotQA).
Concise Rewriting Query noise from redundant or conversational phrases. Improving precision for overly verbose user queries.
Neutralized Claim Reformulation Retriever’s confirmation bias towards a stated claim. Fact-checking and verification (e.g., FEVER, SciFact).
Table 1: Overview of our five expert-crafted rewriting strategies and their targeted applications.
@@ 3.2 Explicit Strategic Primitives
To move past generic prompting, we introduce five explicit, expert-crafted query rewriting strategies, each tailored to address specific challenges in dense retrieval. Rather than relying on simple keyword manipulation, these strategies systematically alter the semantic structure of the query to mitigate common retrieval failures. The primitives, outlined succinctly in Table 1, are not mutually exclusive and often involve inherent trade-offs. For instance, semantic expansion may improve recall at the cost of precision. Collectively, these strategies constitute the discrete action space from which our SAGE framework dynamically selects and applies. The detailed prompts describing each strategy can be found in Appendix A.1.
@@ 3.3 The SAGE Framework
For each input query, the policy model is prompted to generate a batch of $N$ rollouts. A key distinction of our framework is that each action is a structured pair, $\{q_{i},s_{i}\}$, comprising both the rewritten query and the integer ID of the selected strategy. This design compels the LLM not only to generate an effective rewrite but also to commit to an explicit, interpretable strategy, thus making its "intent" transparent.
Each rewritten query $q_{i}$ is then evaluated by the retrieval environment to obtain an initial reward, $r_{\text{orig},i}$, based on its NDCG@10 score. In a conventional RL setup, this raw score would be used directly as the final reward. We posit, however, that this absolute score is an insufficient learning signal, as it provides no direct credit for selecting a high-performing strategy, nor does it effectively distinguish superior rewrites from merely adequate ones within a batch. To address this, the batch of initial rewards is processed by a dedicated reward shaping module (detailed in Section 3.4). This module transforms the raw scores into a more potent learning signal, $r_{\text{final},i}$, which is ultimately used to update the agent’s policy.
@@ 3.5 Countering Reward Hacking with Forced Exploration
In reproducing previous work Jiang et al. (2025); Li et al. (2024), we identify a pervasive form of reward hacking in the query rewriting task. This phenomenon, where an agent exploits loopholes in the reward function to achieve high scores without accomplishing the intended objective, manifests in a particularly deceptive way in our setting. Due to the strong performance of modern retrievers, an agent can often achieve a high baseline reward by simply using the original query. Since the policy model has direct access to this query, it quickly learns that the simplest "hack" is to copy it, securing a high, risk-free score. This behavior traps the agent in a deceptive local optimum, stifling any meaningful exploration of potentially superior rewriting strategies. To directly counteract this, we introduce two mechanisms designed to force exploration.
@@
4 Experiment
4.1 Effectiveness of Rewriting Strategies
To empirically validate the effectiveness of our five expert-crafted strategies (detailed in Section 3.2 and Appendix A.1), we conducted a comprehensive scaling law analysis. Using BGE-en-base-v1.5 Xiao et al. (2023) as the retriever, we compare our Strategy-Guided Prompting (SGP) approach against the strategies proposed by DMQR-RAGLi et al. (2024) using a Best-of-N methodology. The results are presented in Figure 3. Across all four challenging datasets, including HotpotQA Yang et al. (2018), FEVER Thorne et al. (2018), NFCorpus Boteva et al. (2016), and SciFact Wadden et al. (2020), our SGP consistently and substantially outperforms the baseline. This demonstrates that our strategies provide a more effective upper-bound on performance by successfully guiding the LLM towards higher-quality query rewrites. Similar trends of SGP outperforming the baseline are observed when using the Contriever Izacard et al. (2021) retriever, with detailed results provided in Appendix A.2.
Figure 3: Performance scaling laws for query rewriting using the BGE-base-en-v1.5 retriever across four benchmark datasets: HotpotQA, FEVER, NFCorpus, and SciFact. Each subplot compares the upper-bound performance (NDCG@10) of our Strategy-Guided Prompting (SGP) against the strategies from DMQR-RAGLi et al. (2024), evaluated via a Best-of-N methodology across the Qwen3 model series Yang et al. (2025b).
4.2 Comparison with State-of-the-Art Baselines
We benchmark SAGE against a diverse suite of strong baselines to demonstrate its effectiveness. This includes state-of-the-art open-source models Qwen3-4B, Qwen3-8B, Qwen3-14B, GPT-4.1, GPT-o4-mini, Claude-Sonnect-4, Gemini-2.5-Flash, Gemini-2.5-Pro, Deepseek-R1 Yang et al. (2025b); Team et al. (2023); Achiam et al. (2023); Anthropic (2025); Guo et al. (2025). For a fair comparison under our controlled experimental setup, we also re-evaluate prior RL-based methods, including DeepRetrieval Jiang et al. (2025) and the strategies from DMQR-RAG Li et al. (2024). Against these baselines, we report the performance of SAGE enhanced with our two novel reward shaping mechanisms, SAGE-SCS and SAGE-CRS.
This finding has profound practical implications. In a production environment, the long, static portion of our strategic prompt can be pre-processed and its KV cache stored. During inference, the system only needs to generate the short, rewritten query, leading to a dramatic reduction in latency and computational cost. This architectural advantage, which decouples the pre-computable prompt context from the short dynamic generation, makes SAGE a highly attractive solution for real-world, performance-critical retrieval systems.
@@ 4.2 Comparison with State-of-the-Art Baselines
Table 2: SAGE achieves state-of-the-art performance with remarkable efficiency. This table compares our SAGE framework, fine-tuned on Qwen3-4B, against strong baselines including the re-evaluated DeepRetrieval method and significantly larger proprietary models like Gemini-2.5-Pro and GPT-o4mini. All models are evaluated with a maximum response length of 4096 tokens. The results demonstrate that SAGE not only outperforms other RL approaches but also achieves retrieval effectiveness (NDCG@10) competitive with these massive models, while requiring substantially fewer tokens.
5 Ablation Study
5.1 The Impact of Strategic Guidance
To provide a more intuitive understanding, Figure 4 presents a direct case comparison. The output from the model trained without strategies is generic and fails to capture the query’s nuance, whereas the output from SAGE is scientifically precise and far more effective for retrieval.

### v2

@@ Abstract
Query rewriting is pivotal for enhancing dense retrieval, yet current methods demand large-scale supervised data or suffer from inefficient reinforcement learning (RL) exploration. In this work, we first establish that guiding Large Language Models (LLMs) with a concise set of expert-crafted strategies substantially improves retrieval effectiveness on challenging benchmarks, including HotpotQA, FEVER, NFCorpus, and SciFact. Building on this insight, we introduce the Strategy-Adaptive Generation Engine (SAGE), which operationalizes these strategies in an RL framework. SAGE introduces two novel reward shaping mechanisms- Strategic Credit Shaping (SCS) and Contrastive Reward Shaping (CRS)-to deliver more informative learning signals. This strategy-guided approach not only achieves new state-of-the-art NDCG@10 results, but also uncovers a compelling emergent behavior: the agent learns to select optimal strategies, reduces unnecessary exploration, and generates concise rewrites, lowering inference cost without sacrificing performance. Our findings demonstrate that strategy-guided RL, enhanced with nuanced reward shaping, offers a scalable, efficient, and more interpretable paradigm for developing the next generation of robust information retrieval systems.
SAGE: Strategy-Adaptive Generation Engine for Query Rewriting
Teng Wang OPPO Research Institution wt0318@connect.hku.hk          Hailei Gong Tsinghua University gonghl18@mails.tsinghua.edu.cn
Changwang Zhang OPPO Research Institution zhangchangwang@oppo.com          Jun Wang OPPO Research Institution wangjun7@oppo.com
@@ 1 Introduction
Effective information retrieval (IR) is increasingly reliant on dense retrieval systems, which map queries and documents into a shared semantic space. The performance of these systems, however, is fundamentally bound by the quality of the input query. To bridge the significant gap between a user’s initial intent and a query optimized for machine comprehension, query rewriting has emerged as a critical component. While LLMs have shown significant promise for data generation, comprehension, and reasoning Achiam et al. (2023); Wang et al. (2025a); Yang et al. (2025b, a); Wang et al. (2025b), current methodologies face two primary obstacles: traditional supervised fine-tuning demands large-scale, costly manual annotations, whereas modern RL approaches, such as PPO Schulman et al. (2017) and GRPO Shao et al. (2024), often struggle with inefficient exploration. This inefficiency not only hampers the discovery of optimal rewriting strategies but can also result in unstable training dynamics and even catastrophic failures, where models produce incoherent or irrelevant outputs.
@@
Figure 1: An overview of the SAGE framework. SAGE operationalizes expert-crafted strategies within a reinforcement learning loop. The policy model generates a strategy-guided rewrite, which is evaluated against the environment to produce an initial reward (NDCG@10). This reward is then transformed by our novel shaping modules-Strategic Credit Shaping or Contrastive Reward Shaping-to create a potent learning signal for the agent.
@@
The successful application of on-policy RL to fine-tune LLMs is pioneered by algorithms such as PPO Schulman et al. (2017), which in turn inspired a family of variants including GRPO Shao et al. (2024), VAPO Yue et al. (2025), and DAPO Yu et al. (2025). While these powerful algorithms have enabled new capabilities in query rewriting for LLMs, many prior approaches Jiang et al. (2025) simply adopt them as generic optimization tools, without incorporating domain-specific guidance tailored to the query rewriting task.
@@
To bridge this gap, we introduce the Strategy-Adaptive Generation Engine (SAGE), as demonstrated in Fig. 1, a framework that directly integrates our human-designed strategies into the GRPO algorithm, steering LLMs toward more effective query-rewriting policies. To further refine the learning signal, we propose two novel reward shaping schemes. In addition to using the direct NDCG@10 score, we introduce Strategic Credit Shaping (SCS), which assigns credit based on the average performance of each strategy, and Contrastive Reward Shaping (CRS), which transforms absolute scores into relative performance measures.
@@
By reproducing prior approaches Li et al. (2024); Jiang et al. (2025), we identify a significant form of reward hacking Weng (2024); Wang et al. (2024a) inherent in the query rewriting task. When fine-tuned on datasets such as HotpotQA Yang et al. (2018), models tend to converge to the trivial policy of directly copying the input query provided in the prompt, as modern retrievers like BGE-en-v1.5 Xiao et al. (2023) already achieve strong baseline performance with the original queries. Notably, even semantically appropriate rewrites frequently result in lower retrieval scores, discouraging exploration and trapping the agent in a local optimum that precludes the discovery of potentially superior strategies.
@@
To address this issue, we promote exploration by modifying the prompt and penalizing outputs identical to the original query. This approach encourages the agent to venture beyond the safe default. The effectiveness of this strategy is further validated by detailed ablation studies in Section 4.3.
@@
We evaluate SAGE on two challenging datasets: HotpotQA Yang et al. (2018) and NFCorpus Boteva et al. (2016). Our results demonstrate that SAGE achieves state-of-the-art retrieval effectiveness as measured by NDCG@10. Notably, we observe an emergent behavior where SAGE learns a more efficient reasoning process, substantially reducing token usage. This improved efficiency directly translates to lower inference latency and reduced computational costs.
@@
    We show that prompting LLMs with a small set of interpretable strategies substantially improves query rewriting quality, establishing the performance upper bound attainable through prompting alone.
@@
    We introduce SAGE, a novel RL framework that systematically integrates explicit strategies into the learning process. SAGE autonomously adapts strategy selection via our proposed reward-shaping mechanisms, enabling more effective policy optimization.
@@
    We establish a new state-of-the-art in dense retrieval effectiveness (NDCG@10) using SAGE, while identifying a notable emergent behavior: SAGE learns a more efficient reasoning process, substantially reducing inference latency and computational costs.
@@ 2 Methodology
The core idea of our methodology is to replace black-box RL optimization with explicit, human-interpretable decision-making. To this end, we propose the Strategy-Adaptive Generation Engine (SAGE), which reformulates query rewriting as a structured, strategy-driven decision process. Our framework directly incorporates expert-crafted strategies into RL training to systematically guide both exploration and policy learning, as illustrated in Figure 1.
@@ 2.1 Problem Formulation
We frame query rewriting as an RL task. Given an initial user query $q_{orig}$ and a document collection $\mathcal{D}$, the objective is to learn a policy $\pi$ that generates an improved query $q$. The environment evaluates the rewritten query using NDCG@10, yielding an initial reward $r_{\text{orig}}=\text{NDCG@10}(q,\mathcal{D})$, which is further refined by our reward shaping methods (see Section 2.4) to produce the final reward $r_{\text{final}}$. The agent thus aims to maximize the expected final reward: $\mathbb{E}_{q\sim\pi(\cdot|q_{orig})}[r_{\text{final}}]$. We use GRPO Shao et al. (2024) as our optimization algorithm and incorporate a set of expert-crafted rewriting strategies $\mathcal{S}=\{s_{1},s_{2},\dots,s_{5}\}$, providing explicit guidance for query reformulation.
@@ 2.2 Explicit Strategic Primitives
To move beyond generic prompting, we introduce five explicit, expert-crafted query rewriting strategies, each tailored to address specific challenges in dense retrieval. These strategies systematically alter the semantic structure of queries to mitigate common retrieval failures. The primitives, summarized in Table 4, are not mutually exclusive and often involve inherent trade-offs (e.g., semantic expansion may improve recall at the expense of precision). Collectively, they constitute the discrete action space from which our SAGE framework dynamically selects and applies. Detailed prompt templates for each strategy are provided in Appendix A.1.
@@ 2.3 The SAGE Framework
For each input query, the policy model generates a batch of $N$ rollouts, where each action is a structured pair, $\{q_{i},s_{i}\}$, comprising the rewritten query and the integer ID of the selected strategy. This formulation requires the LLM to produce not only an effective rewrite but also an explicit strategy choice, rendering its intent interpretable.
Each rewritten query $q_{i}$ is evaluated by the retrieval environment to obtain an initial reward $r_{\text{orig},i}$, measured by NDCG@10. In conventional RL, this raw score serves as the final reward; however, we argue that it provides insufficient learning signal, as it neither credits effective strategy selection nor clearly distinguishes superior rewrites within a batch. To address this, we process the initial rewards with a dedicated reward shaping module (see Section 2.4), which transforms the raw scores into a more informative signal $r_{\text{final},i}$ for policy updates.
@@ 2.5 Countering Reward Hacking with Forced Exploration
Reproducing previous work Jiang et al. (2025); Li et al. (2024), we observe pervasive reward hacking in query rewriting: agents exploit reward function loopholes to achieve high scores without fulfilling the intended objective. In our setting, the strong baseline performance of modern retrievers allows agents to obtain high rewards simply by copying the original query, a strategy the policy model quickly adopts since it has direct access to the input. As a result, the agent becomes trapped in a deceptive local optimum, stifling exploration of potentially superior rewriting strategies. To directly counteract this, we introduce two mechanisms designed to force exploration.
@@
Figure 3: Performance scaling laws for query rewriting using the BGE-base-en-v1.5 retriever across four benchmark datasets: HotpotQA, FEVER, NFCorpus, and SciFact. Each subplot compares the upper-bound performance (NDCG@10) of our Strategy-Guided Prompting (SGP) against the strategies from DMQR-RAGLi et al. (2024), evaluated via a Best-of-N methodology across the Qwen3 model series Yang et al. (2025a).
@@
Table 1: SAGE achieves state-of-the-art performance with remarkable efficiency. This table compares our SAGE framework, fine-tuned on Qwen3-4B, against strong baselines including the re-evaluated DeepRetrieval method and significantly larger proprietary models like Gemini-2.5-Pro and GPT-o4-mini. All models are evaluated with a maximum response length of 4096 tokens. The results demonstrate that SAGE not only outperforms other RL approaches but also achieves retrieval effectiveness (NDCG@10) competitive with these massive models, while requiring substantially fewer tokens.
3 Experiment
3.1 Effectiveness of Rewriting Strategies
To empirically validate our five expert-crafted strategies (detailed in Section 2.2 and Appendix A.1), we conduct a comprehensive scaling law analysis using BGE-en-base-v1.5 Xiao et al. (2023) as the retriever. We compare our Strategy-Guided Prompting (SGP) to the strategies in DMQR-RAG Li et al. (2024) under a Best-of-N methodology, with results shown in Figure 3. Across all four challenging datasets, including HotpotQA Yang et al. (2018), FEVER Thorne et al. (2018), NFCorpus Boteva et al. (2016), and SciFact Wadden et al. (2020), our SGP consistently and substantially outperforms the baseline. This demonstrates that our strategies provide a more effective upper-bound on performance by successfully guiding the LLM towards higher-quality query rewrites. Similar trends of SGP outperforming the baseline are observed when using the Contriever Izacard et al. (2021) retriever, with detailed results provided in Appendix A.2.
3.2 Comparison with State-of-the-Art Baselines
Unless otherwise specified, all experiments are conducted using BGE-en-base-v1.5. Detailed training settings are provided in Appendix A.3.
We benchmark SAGE against a diverse suite of strong baselines to demonstrate its effectiveness. This includes state-of-the-art open-source models Qwen3-4B, Qwen3-8B, Qwen3-14B, GPT-4.1, GPT-o4-mini, Claude-Sonnect-4, Gemini-2.5-Flash, Gemini-2.5-Pro, Deepseek-R1 Yang et al. (2025a); Team et al. (2023); Achiam et al. (2023); Anthropic (2025); Guo et al. (2025). For a fair comparison under our controlled experimental setup, we also re-evaluate prior RL-based methods, including DeepRetrieval Jiang et al. (2025) and the strategies from DMQR-RAG Li et al. (2024). Against these baselines, we report the performance of SAGE enhanced with our two novel reward shaping mechanisms, SAGE-SCS and SAGE-CRS.
This finding has profound practical implications. In a production environment, the long, static portion of our strategic prompt can be pre-processed and its KV cache stored. At inference time, the system only needs to generate a short reasoning, significantly reducing latency and computational cost. This architectural advantage, which decouples the pre-computable prompt context from the short dynamic generation, makes SAGE a highly attractive solution for real-world, performance-critical retrieval systems.